An investor owns $5,000 of Adobe System stock $6,000 of Dow Chemical and $9,000 of Office Depot. What are the portfolio weights of each stock?

Adobe System
Dow Chemical
Office Depot
Answer #1

Total value=(5000+6000+9000)=$20,000

Portfolio weight
Adobe System (5000/20,000)=25%(or 0.25)
Dow Chemical (6000/20,000)=30%(or 0.30)
Office Depot (9000/20,000)=45%(or 0.45)